Earth soil is being delivered to an embankment at a rate of 2000 lcy/hr. The compaction process is being done using rollers with a drum width of 4.16 ft. The recommended compacted lift height is 7 inches with 4 drum passes at a speed of 2 mph. With an assumption of 55 min/hour efficiency, determine the number of rollers required to provide the necessary compaction for the delivered soil.


Assume that lifts are placed at 10-in thickness before compaction for your
calculations.
Answer #1

Here we have,

Required delivery rate of earth soil for embankment = 2000 lcy/hr

Width of roller ,W = 4.16 ft

Compacted lift height,L = 7 inches

Number of passes,N = 4

Speed ,S = 2 mph

Efficiency,η = 55 min/ hr = 91.67%

Lift thickness before compaction = 10 inches

Now as we know that production rate of roller is given by,

\small R = \frac{16.3*W*S*L}{N}*\eta

Where,

R = Production rate in Compacted Cubic yards per hour, i.e ccy/hr

16.3 = Factor for unit conversion.

W = width of roller in feet,ft

S = Speed of vehicle in miles per hour,mph

L = Compacted lift thickness in inches,in

N = Number of passes.

η = Efficiency

Now the values given in the formula,we get

\small R = \frac{16.3*4.16*2*7}{4}*0.9167

\small R = 217.55 \ ccy/hr

Now as the delivery rate is given in loose cubic yards per hour,i.e lcy/hr,we have to convert this into ccy/hr.

As the thickness of lift before compaction was 10 inches and after the compaction it is 7 inches,

Lcy to ccy conversion factor is (7/10) = 0.7.

Thus required dilivery rate in ccy/hr is 2000 × 0.7 ccy/hr = 1400 ccy/hr.

Minimum Number of rollers required is given by,

\small =\frac{\ Required\ delivery\ rate}{\ Production\ rate\ of\ roller}=\frac{1400 \ ccy/hr}{217.55 \ ccy/hr}=6.43

Thus atleast 7 Rollers required to deliver 2000 lcy/hr of earth soil.